# Integral modelling and Reinforcement Learning control of 3D liquid metal coating on a moving substrate

### v1

@@ VI Reinforcement Learning Control
The undulation control problem consists in finding the optimal weights $\mathbf{w}^{*}$ of a parameterized feedback control law $\pi_{\mathbf{w}}(\mathbf{s})$, which minimizes an objective functional accounting for the amplitude of undulation waves. Based on $n_{s}$ observations of the liquid film thickness $\mathbf{s}=\mathbf{s}(\hat{h}(\hat{t};\mathbf{w}))\in\mathbb{R}^{n_{\rm s}}$ sampled at time $\hat{t}$, the parametric feedback control law $\pi_{\mathbf{w}}(\mathbf{s})$ gives the value of control actions $\mathbf{a}=\mathbf{a}(\hat{t})\in\mathbb{R}^{n_{\rm a}}$, which corresponds to the nozzle exit velocities $U_{j}$ or magnetic field intensities $b_{j}$ of the $n_{\rm a}$ actuators. The objective functional to minimize is given by the time integral of a reward function $r(\hat{t};\mathbf{w})$ over a liquid film simulation that lasts for a duration $T$. In this setting, the undulation control problem is expressed as:
[…]
@@
$\mathcal{L}(\hat{h}(\hat{t};\mathbf{w}))=\exp\left(\text{std}\left(50(\hat{h}(\hat{t};\mathbf{w})-\hat{h}_{0})\right)\right)-1.$ (38)
where $\hat{h}_{0}$ represents the initial thickness of the flat film at $\hat{t}=0$, exp denotes the exponential function and std denotes the standard deviation computation in both space and time.
@@ VI.1 Proximal Policy Optimization
The undulation control problem (36) with (37) and (38), constrained by the 3D IBL equations (22), is solved using the Proximal Policy Optimization (PPO) reinforcement learning algorithm [27] using the implementation provided by the Stable Baselines 3 library [47].
This method uses a stochastic feedback control law, with actions $\mathbf{a}$ sampled from a parameterized multivariate Gaussian distribution $\pi_{\mathbf{w}}(\mathbf{a}|\mathbf{s})=\mathcal{N}(\mu_{\mathbf{w}}(\mathbf{s}),\sigma_{\mathbf{w}}(\mathbf{s}))$. The vectors $\mu_{\mathbf{w}}(\mathbf{s})\in\mathbb{R}^{n_{\rm a}}$ and $\sigma_{\mathbf{w}}(\mathbf{s})\in\mathbb{R}^{n_{\rm a}}$ represent the mean and standard deviation of the distribution, and the symbol |||| indicates the conditioning of observations $\mathbf{s}$. For any given set of observations $\mathbf{s}$, $\mu_{\mathbf{w}}(\mathbf{s})$ and $\sigma_{\mathbf{w}}(\mathbf{s})$ are provided by a neural network with weights $\mathbf{w}$ of size $256\times 256$ neurones. Using a stochastic feedback control law, variability is introduced into the action selection process, fostering the exploration of different control strategies, which ultimately enhances control performance.
@@
The PPO seeks the optimal parametrization $\mathbf{w}^{*}$ through a trial-and-error process, which consists of running 300 liquid film simulations in parallel, evaluating the instantaneous performance of the feedback control law, learning from the accumulated experience, and then updating its parameterization in a gradient descent fashion.
@@
Figure 5: Scheme of the reinforcement learning process where actions $\mathbf{a}k$ are selected according to the stochastic control law $\pi$, and then passed to the numerical environments. These environments integrate the IBL equations and return liquid film thickness observations $\mathbf{s}{k+1}$ and rewards $r_{k+1}$. Based on the accumulated knowledge stored in the buffer, the control law parameters are updated.
@@
Figure 5 shows the main steps of this process: actions are sampled from the Gaussian distribution $\pi_{\mathbf{w}_{n}}(\mathbf{a}_{k}|\mathbf{s}_{k-1})$, parameterized by the neural network weights $\mathbf{w}_{n}$ (white box). These actions are then passed on to $N_{e}=30$ liquid film simulations, known as numerical environments, at time $\hat{t}_{k}=(k\,m\,-1)\Delta\hat{t}$, where $\Delta\hat{t}$ represents the integration time step. To accelerate the optimization process, these numerical environments are executed in parallel across multiple CPUs (grey boxes), simulating shifted undulation waves to increase variability and enhance the chances of learning a better parameterization (more details are provided in Subsection VI.2).
@@
The actions $\mathbf{a}_{k}$, which differ across environments, are held constant for $m=30$ integration time steps to prevent sudden changes that could hamper the PPO learning process. At the end of this period, each numerical environment outputs a vector of liquid film observations $\mathbf{s}_{k+1}=\mathbf{s}(\hat{h}(\hat{t}_{k+1};\mathbf{w}_{n}))$ evaluated at the last timestep of the integration period $\hat{t}_{k+1}$, along with a total reward function $R_{k+1}$ given by the sum of the reward function at each time step, which reads:
$R_{k+1}=\sum_{j=1}^{m}r(\hat{t}=(k\,m+j-1)\Delta\hat{t};\mathbf{w}_{n})).$ (39)
@@
The sequence of actions, observations, and rewards gathered during a simulation of duration $\hat{t}=T$ constitutes a trajectory. The final time $T$ is defined as $T=\Delta\hat{t}(K,m-1)$, where $K$ represents the algorithm’s total number of interactions with the numerical environment per simulation. Trajectories from different simulations and numerical environments are stored in a buffer (blue box).
[…]
@@
where $\gamma\in[0,1]$ is the discount factor that determines the weight of future rewards, and $V(\mathbf{s}_{k})$ represents the value function. This function calculates the expected sum of rewards starting from a given state $\mathbf{s}_{k}$, with actions chosen according to the control law $\pi_{\mathbf{w}_{n}}$ until the last interaction with the numerical environments ($k=K$) and it is defined as follows:
$V(\mathbf{s}_{k})=\mathbb{E}\left[\sum_{j=k}^{K}\gamma^{j-k}r_{j}\right]$ (42)
The value function is approximated by a separate neural network of size $256\times 256$ neurones and is updated using the sum of the rewards collected in the buffer.
@@
$g(\iota,A)=\begin{cases}(1+\iota)A&\text{if }A\leq 0,\\ (1-\iota)A&\text{if }A>0.\end{cases}$ (43)
@@ VI.2 Control test cases
This section describes the numerical environment used to solve the control problem introduced in Section VI using the PPO algorithm presented in Subsection.
@@
Figure 6: Scheme of the spectral environment with the physical domain $\Omega_{1}$ containing the actuators (black crosses), the observations (red dots), the forcing jets (blue circles) and the reward area (light orange region), surrounded by an absorbing perfectly matched layer (green region) with periodic boundary conditions on the found boundaries ($\partial\Omega$).
@@
where $\omega$ is the angular frequency of the liquid film waves and $\sigma_{x}(\hat{x},\hat{z})$ and $\sigma_{z}(\hat{x},\hat{z})$ are positive functions which determine the strength of the absorbing layer, defined as:
@@
The transformation in equation (45) involves introducing a set of auxiliary equations, discretized both in space and time, similar to the IBL equations. For convenience, these equations are provided in Appendix C.
Moving to the undulation instability, this appears in the liquid film as two-dimensional waves propagating streamwise, with minimal variations along $\hat{z}$. Numerical simulations from Large Eddies Simulations (LES) [51] and 2D IBL model [17] indicate that these waves are characterized by a nondimensional frequency in the range $[0.05,0.2]$, a peak-to-peak amplitude ranging from $5\%$ to $10\%$ of the unperturbed film thickness, and wavelengths within the range $\lambda\in[25,35]\,\text{mm}$.
To reproduce these characteristics, we use a set of blowing and sucking gas jets. The nozzle exit velocity of these jets $U_{j,f}^{i}$ varies with a harmonic function of the nondimensional time $\hat{t}$, with a random phase shift $\phi\in[0.2\pi,0.5\pi]$ to induce a small variation in the spanwise direction. The exit velocity of the $j$-th forcing jets reads:
$U_{j,f}^{i}=\grave{A}_{fj}\cos{(2\pi\hat{f}_{fj}\hat{t}+\phi_{i})}.$ (47)
To introduce greater variability into the simulation and promote the learning of robust control functions, the amplitude $\grave{A}_{fj}$ and the frequency $\hat{f}_{fj}$ are modelled as finite Fourier series $y_{fj}$ with randomly selected coefficients $a_{j}$ and $b_{j}$ [52], reading:
$\begin{split}y_{fj}(\hat{t})=\sqrt{2}\sum_{j=1}^{4}\Big{[}a_{j}\cos{\left(2\pi j\hat{t}\right)+b_{j}\sin{(2\pi j\hat{t})}}\Big{]},\end{split}$ (48)
@@
[…]
Figure 7: 3D plot of the non-dimensional liquid film showing the preparation of the undulation instability waves generated by the forcing gas jets.
@@
[…]
@@
where $(U_{j})_{i+1,e}$ is the action passed to the solver, $(U_{j})_{i+1,e}^{ML}$ is the action selected by the PPO and $(U_{j})_{i-1,e}$ is the action passed to the solver at the previous interaction with the film.
Three distinct actuator configurations were analysed: a standalone 2D gas jet, a standalone 2D electromagnet, and a combination of both actuators. A harmonic control action was applied with amplitude, frequency, and phase variations throughout the simulation in the standalone configurations.
Figure 8 illustrates the numbering and placement of the three liquid film observations (red dashed lines) and the actuators (black dashed and dash-dotted lines) for the case with a single (left) and double (rights) actuators.
Figure 8: Schematic showing the numbering of observations and actuators in the 3D undulation control environment with zinc.
The three liquid film observations are the average liquid film thickness sampled at four points along $\hat{z}$ at three locations along $\hat{z}$. For a single 2D actuator, these are placed at $\hat{x}=(32,31,30)(x=(5.16,5,4.83)[\rm{cm}]]$ and $\hat{z}=(6.3,9.4,12.56,15.7)((1,1.5,2,2.5)\,\rm{cm})$. For the two 2D actuators, these are placed at $\hat{x}=(33,31,29)((5.32,5,4.68)\,\rm{cm})$ and $\hat{z}=(6.2,9.4,12.6,15.8)((1,1.5,2,2.5)\,\rm{cm})$.
@@ VII Results
This section is divided into three subsections. Subsection VII.1 reports on the asymptotic behaviour of the steady-state solution and the closure terms for a magnetic field approaching zero. Subsection  VII.2 reports on the wave dynamics that arises when setting the non-linear phase speed in the leading order in the IBL model. Finally, subsection  VII.3 reports on the results of Reinforcement Learning control of the unstable undulation perturbations using gas jets and electromagnetic actuators, considering both standalone and tandem configurations.

### v2

@@ VI Reinforcement Learning Control
The undulation control problem consists of finding the optimal weights $\mathbf{w}^{*}$ of a parametrised feedback control law $\pi_{\mathbf{w}}(\mathbf{s})$, which minimises an objective functional accounting for the amplitude of undulation waves. Based on $n_{s}$ observations of the liquid film thickness $\mathbf{s}=\mathbf{s}(\hat{h}(\hat{t};\mathbf{w}))\in\mathbb{R}^{n_{\rm s}}$ sampled at time $\hat{t}$, the parametric feedback control law $\pi_{\mathbf{w}}(\mathbf{s})$ gives the value of control actions $\mathbf{a}=\mathbf{a}(\hat{t})\in\mathbb{R}^{n_{\rm a}}$, which corresponds to the nozzle exit velocities $U_{j}$ or magnetic field intensities $b_{j}$ of the $n_{\rm a}$ actuators. The objective functional to minimise is given by the time integral of a reward function $r(\hat{t};\mathbf{w})$ over a liquid film simulation that lasts for a duration $T$. In this setting, the undulation control problem is expressed as:
$\min_{\mathbf{w}}\int_{0}^{T}r(\hat{t};\mathbf{w})\,d\hat{t}$. (42)
@@
$\mathcal{L}(\hat{h}(\hat{t};\mathbf{w}))=\exp\left(\text{std}\left(50(\hat{h}(\hat{t};\mathbf{w})-\hat{h}_{0})\right)\right)-1,$ (44)
where $\hat{h}_{0}$ represents the initial thickness of the flat film at $\hat{t}=0$, expexp denotes the exponential function and stdstd denotes the standard deviation.
@@ VI.1 Proximal Policy Optimization
The undulation control problem (42) with (43) and (44), constrained by the 3D IBL equations (28), is solved using the Proximal Policy Optimisation (PPO) reinforcement learning algorithm [45] using the implementation provided by the Stable Baselines 3 library [43].
The PPO algorithm selects actions according to a stochastic control policy whose parameters are progressively updated during training. Given a set of liquid film observations $\mathbf{s}$, the control actions $\mathbf{a}$, representing the nozzle exit velocity and/or the magnetic field intensity, are sampled from a parameterised multivariate Gaussian distribution $\mathcal{N}$, reading:
$\mathbf{a}\sim\mathcal{N}\big(\mu_{\mathbf{w}}(\mathbf{s}),\,\sigma_{\mathbf{w}}(\mathbf{s})\big),$ (45)
where $\mu_{\mathbf{w}}(\mathbf{s})$ is the mean and $\sigma_{\mathbf{w}}(\mathbf{s})$ is the standard deviation. The value of $\mu_{\mathbf{w}}(\mathbf{s})$ and $\sigma_{\mathbf{w}}(\mathbf{s})$ is given by the function $\pi_{\mathbf{w}}(\mathbf{a}|\mathbf{s})$, which is represented by a neural network with weights $\mathbf{w}$ which consists of two hidden layers of 256 neurons each. This stochastic feedback control formulation introduces variability into the action-selection process, thereby encouraging exploration of alternative control strategies and improving overall control performance.
@@
The PPO seeks the optimal parametrisation $\mathbf{w}^{*}$ through a trial-and-error process that runs 300 liquid film simulations in parallel, evaluates the instantaneous performance of the feedback control law, learns from accumulated experience, and then updates the parameterisation via gradient descent.
@@
Figure 2: Schematic representation of the reinforcement learning process, in which actions $\mathbf{a}_{k}$ are selected according to the stochastic control policy $\pi_{\mathbf{w}_{n}}$ and passed to the numerical environments (grey boxes). These environments integrate the IBL equations and give updated liquid film thickness observations $\mathbf{s}_{k+1}$ and corresponding rewards $r_{k+1}$. The control policy parameters are then updated based on the rewards and observations stored in the buffer (blue box).
@@
Figure 2 shows the main steps of this process: actions are sampled from the Gaussian distribution $\pi_{\mathbf{w}_{n}}(\mathbf{a}_{k}|\mathbf{s}_{k-1})$, parameterized by the neural network weights $\mathbf{w}_{n}$ (white box). These actions are then passed on to $N_{e}=30$ liquid film simulations, known as numerical environments, at time $\hat{t}_{k}=(k\,m\,-1)\Delta\hat{t}$, where $\Delta\hat{t}$ represents the integration time step. To accelerate the optimisation process, these numerical environments are executed in parallel across multiple CPUs (grey boxes), simulating shifted undulation waves to increase variability and improve the likelihood of learning a better parameterisation (see Subsection VI.2 for details).
@@
The actions $\mathbf{a}_{k}$, which differ across environments, are held constant for $m=30$ integration time steps to prevent sudden changes that could hamper PPO learning. At the end of this period, each numerical environment outputs a vector of liquid film observations $\mathbf{s}_{k+1}=\mathbf{s}(\hat{h}(\hat{t}_{k+1};\mathbf{w}_{n}))$ evaluated at the last timestep of the integration period $\hat{t}_{k+1}$, along with a total reward function $R_{k+1}$ given by the sum of the reward function at each time step, which reads:
$R_{k+1}=\sum_{j=1}^{m}r(\hat{t}=(k\,m+j-1)\Delta\hat{t};\mathbf{w}_{n}).$ (46)
@@
The sequence of actions, observations, and rewards gathered during a simulation of duration $\hat{t}=T$ constitutes a trajectory. The final time $T$ is defined as $T=\Delta\hat{t}(K,m-1)$, where $K$ represents the algorithm’s total number of interactions with the numerical environment per simulation. Trajectories from different simulations and numerical environments are stored in a buffer (blue box in Fig. 2).
Leveraging the knowledge of liquid film dynamics acquired from stored trajectories, the PPO algorithm enhances control performance by updating the weights $\mathbf{w}_{n}$ every time $(k\cdot N_{e}\;\rm{mod}\;2048=1)$. This update is made by maximizing a clipped objective function, which entails the minimum of the probability ratio between the updated control law $\pi_{\mathbf{w}_{n+1}}(\mathbf{a}|\mathbf{s})$ and the old control law $\pi_{\mathbf{w}_{n}}(\mathbf{a}|\mathbf{s})$, along with a clipping function $g=g(\iota,A)$ [39]. The maximisation problem is formulated as:
@@
where $\gamma\in[0,1]$ is the discount factor that determines the weight of future rewards, and $V(\mathbf{s}_{k})$ represents the value function. This function calculates the expected sum of rewards starting from a given state $\mathbf{s}_{k}$, with actions chosen according to the control law $\pi_{\mathbf{w}_{n}}$ until the last interaction with the numerical environments ($k=K$), and it is defined as follows:
$V(\mathbf{s}_{k})=\mathbb{E}\left[\sum_{j=k}^{K}\gamma^{j-k}r_{j}\right].$ (49)
The value function is approximated by a separate neural network with $256\times 256$ neurons and is updated using the sum of rewards collected in the buffer.
@@
$g(\iota,A)=\begin{cases}(1+\iota)A&\text{if }A\leq 0,\\ (1-\iota)A&\text{if }A>0,\end{cases}$ (50)
@@ VI.2 Control test cases
This section describes the numerical environment used to solve the control problem introduced in Section VI using the PPO algorithm presented in Subsection VI.1.
@@
Given that the liquid film thickness is of order $O(10^{-1})$, the nondimensional group $\Lambda=9.2$, $\hat{b}\in[0,1]$ (see Table 1), and the spatial derivatives of the film thickness are small, the Maxwell stress terms in the normal (26b) and tangential (26c) boundary conditions can be regarded as small quantities and are therefore neglected from our calculations for simplicity.
@@
Figure 3: Rectangular domain with periodic boundary conditions used in the undulation control problem, showing the positions of the control jets (red crosses) and control magnets (blue squares), the observation points (red dots), the forcing jets (blue circles), and the reward area (light orange region), with the perfectly matched layer (green region).
@@
[…]
@@
The transformation in equation (52) involves introducing a set of auxiliary equations, discretised both in space and time, similar to the IBL equations. For convenience, these equations are provided in Appendix C.
Moving to undulation instability, this manifests in the liquid film as two-dimensional waves propagating streamwise, with minimal variation along $\hat{z}$. Numerical simulations from Large Eddies Simulations (LES) [3, 1] and 2D IBL model [32] indicate that these waves are characterized by a nondimensional frequency in the range $[0.05,0.2]$, a peak-to-peak amplitude ranging from $5\%$ to $10\%$ of the unperturbed film thickness, and wavelengths within the range $\lambda\in[25,35]\,\text{mm}$.
These waves originate from the interaction between the wiping gas jet and the liquid film [1]. To emulate comparable undulation patterns over short distances, we introduce an artificial forcing mechanism that consists of numerical jets represented as source terms in the governing equations. These sources inject alternating positive and negative perturbations into the local flow rate, effectively seeding waves with amplitudes and wavelengths consistent with those reported in [3, 1]. The formulation of these sources follows the correlations defined for the control actuators (see Sec. IV), but allows negative input velocities. Although such negative forcing is non-physical, the goal here is not to represent realistic jet actuation but rather to generate synthetic disturbances that reproduce the characteristic undulation instability within the computationally accessible domain. For these forcing jets, the velocity of the $j$-th jet is defined as
$U_{j,f}^{i}=\grave{A}_{fj}\cos(2\pi\hat{f}_{fj}\hat{t}+\phi_{i}),$ (54)
where $\phi_{i}\in[0.2\pi,0.5\pi]$ introduces mild spanwise variability. Moreover, to introduce greater variability into the simulation and promote the learning of robust control functions, the amplitude $\grave{A}_{fj}$ and the frequency $\hat{f}_{fj}$ are modelled as finite Fourier series $y_{fj}$ with randomly selected coefficients $a_{j}$ and $b_{j}$ [16], reading:
[…]
@@
Figure 4 shows the nondimensional uncontrolled liquid film with the undulation instability induced by the forcing jets propagating through the domain. The wavelength ranges from $25\,\text{mm}$ to $27\,\text{mm}$, in the range of most amplified waves identified in [32].
Figure 4: 3D plot of the non-dimensional liquid film showing the propagation of the undulation instability waves generated by the forcing gas jets.
@@
The pressure and shear distributions of the control and forcing jets are modelled using experimental correlations for 3D impinging circular jets (see Subsection IV), with the velocity at the jet nozzle exit, $U_{j}\in[0,50]\,\text{m/s}$, used as the control parameter. To avoid sharp changes in the control action, this is smoothed with an exponential moving average, reading:
@@
where $(U_{j})_{i+1,e}$ is the action passed to the solver, $(U_{j})_{i+1,e}^{ML}$ is the action selected by the PPO, and $(U_{j})_{i-1,e}$ is the action passed to the solver at the previous interaction with the film.
Three actuator configurations were analysed: a standalone 2D gas jet, a standalone 2D electromagnet, and a combination of both actuators. For these, two control policies were considered. The first, referred to as the harmonic policy, prescribes the actuation signal in a sinusoidal form,
$\hat{u}_{c}(t_{k})=\acute{A}(\mathbf{s}_{k})\,\sin\!\big(2\pi\acute{f}(\mathbf{s}_{k})\,t+\acute{\phi}(\mathbf{s}_{k})\big),$
where $\hat{u}_{c}$ denotes the actuator command, corresponding to either the jet velocity $U_{j}$ or the magnetic field intensity $\hat{b}_{t}$, and $\mathbf{s}_{k}$ is the state vector containing the observed film-height information at time step $k$. The parameters $\acute{A}(\mathbf{s}_{k})$, $\acute{f}(\mathbf{s}_{k})$, and $\acute{\phi}(\mathbf{s}_{k})$ are determined by the PPO policy as functions of the instantaneous system state. This formulation, applied only to single–actuator configurations (either a jet or an electromagnet), was introduced to bias the learning process toward harmonic actuation while still allowing the policy to decide whether the forcing should remain strictly periodic. Suppose the learned parameters converge to nearly constant values. In that case, the resulting actuation corresponds to a classical phase–opposition control law, where the actuator oscillates at the dominant wave frequency but in opposite phase to attenuate the free–surface deformation. If, instead, the policy continuously modulates $\acute{A}$, $\acute{f}$, or $\acute{\phi}$, the control signal departs from a pure sinusoid and adapts its waveform to mitigate the incoming, not perfectly periodic, wave pattern.
In the second formulation, referred to as the non-harmonic policy, the PPO agent directly outputs the actuator command at each time step as a function of the instantaneous film-height observations. This corresponds to a fully closed-loop feedback control law in which the policy learns a mapping from the system state to the continuous actuation signal. The non-harmonic policy, therefore, represents a generalisation of the previous case, enabling non-periodic and adaptive control strategies that can respond to arbitrary disturbances in real time.
Figure 5 illustrates the numbering and placement of the three liquid film observations (red dashed lines) and the actuators (black dashed lines) for the case with a single (left) and double (right) actuators.
The three liquid film observations are given by the average of the liquid film thickness sampled at four points along $\hat{z}$ at three locations along $\hat{x}$. For a single 2D actuator, these are placed at $\hat{x}=(32,31,30)(x=(5.16,5,4.83)\rm{cm}$ and $\hat{z}=(6.3,9.4,12.56,15.7)((1,1.5,2,2.5)\,\rm{cm})$. For the two 2D actuators, these are placed at $\hat{x}=(33,31,29)((5.32,5,4.68)\,\rm{cm})$ and $\hat{z}=(6.2,9.4,12.6,15.8)((1,1.5,2,2.5)\,\rm{cm})$.
Figure 5: Schematic showing the numbering of observations and actuators in the 3D undulation control environment.
@@ VII Results
This section is divided into three subsections. Subsection VII.1 reports on the asymptotic behaviour of the steady-state solution and the closure terms for a magnetic field approaching zero. Subsection  VII.2 reports on the wave dynamics that arise when setting the non-linear phase speed in the leading order in the IBL model. Finally, subsection  VII.3 reports on the results of Reinforcement Learning control of the unstable undulation perturbations using gas jets and electromagnetic actuators, considering both standalone and tandem configurations.